# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@ 2.3. Feature Learning Stage
The Feature Learning Stage is fundamental to improving the adaptability and robustness of Wi-Fi sensing systems across diverse domains. This stage includes several key techniques aimed at overcoming the challenges of domain variability, limited labeled data, and environmental changes. Methods such as domain alignment, reconstruction structures, embedding learning, data augmentation, pseudo-labeling and pure feature learning work together to enhance model performance, ensuring that features learned from one domain can be successfully applied to others. These approaches through techniques like adversarial training, similarity metrics, data synthesis, and embedding refinement, enable Wi-Fi sensing models to generalize effectively in dynamic and heterogeneous real-world environments.
2.3.1. Domain Alignment
Domain alignment techniques are crucial for reducing the gap between source and target domains in Wi-Fi sensing systems. These methods ensure better generalization by aligning feature distributions, enabling models to adapt to new environments. In the domain alignment with domain discriminator approach, JADA (Zou et al., 2018a) applies domain adversarial training to classify features as belonging to either the source or target domain, facilitating robust feature alignment. CrossGR (Li et al., 2021b) uses a GAN to generate synthetic data and employs a discriminator to distinguish between real and synthetic samples, ensuring that domain alignment is achieved effectively. Similarly, WiHARAN (Wang et al., 2021a) incorporates a domain discriminator to align activity features across different environments, thus enhancing activity recognition. In a similar vein, i-Sample (Zhou et al., 2024) augments datasets through gradient learning while utilizing a discriminator to identify and align both real and synthetic samples, improving the adaptability of the model. Zhang et al. (Zhang et al., 2021b) employs a feature encoding network, which is deployed in the local server of each environment. The Johnson-Lindenstrauss transform is applied to preserve privacy information before it is sent to the central server. An environment classifier distinguishes which environment the input datum comes from, and as the system converges, it becomes environment-invariant, making it incapable of distinguishing between environments.
2.3.2. Adversarial Domain Adaptation
Adversarial Domain Adaptation (Ganin et al., 2016; Tzeng et al., 2017) aims to align the source and target domains with adversarial training, enabling Wi-Fi sensing to generalize across new individuals, environments, and other novel domains. For example, JADA (Zou et al., 2018a) feeds source-domain data into a source-domain encoder and target-domain data into a target-domain encoder. Through adversarial training with a domain discriminator, the features extracted by both encoders become indistinguishable regarding their domain of origin. The source-domain encoder is then frozen, and a classification head is trained using labeled source-domain data for action recognition. Finally, the target-domain encoder and classification head are combined for inference on target-domain data, achieving cross-domain action recognition. CrossGR (Li et al., 2021b) trains an eliminator to make action-related features unclassifiable, isolating components irrelevant to the actions. By removing these irrelevant components, Wi-Fi features are better suited for cross-person and cross-environment action recognition. In Person-in-WiFi (Wang et al., 2019f), pairs of data inputs are used to pre-train a discriminator that determines whether the data pairs come from the same environment. A generator is then trained to ensure that the features of these pairs are consistently judged as originating from the same environment. This process removes environmental information and enhances the accuracy of cross-scene human pose estimation.
Domain Alignment with Domain Classifier approaches explicitly classify features into domains and use this feedback for alignment. EI (Jiang et al., 2018c) employs a domain classifier with the objective of minimizing the loss function to maximize the performance of domain label prediction, which contrasts with the ultimate goal of learning domain-independent features for activities. CsiGAN (Xiao et al., 2019), supported by CycleGAN (Zhu et al., 2017a), generates virtual samples and employs a domain classifier to distinguish real and synthetic data for cross-domain adaptation. Wi-Adaptor (Zhang et al., 2021e) uses domain classifiers and adversarial training to align features across source and target domains. GESFI (Zhang et al., 2024a) and Khattak (Khattak and Khan, 2022) incorporate domain classifiers to address variability in gesture recognition and location-based sensing, respectively. DINN (Zhou et al., 2021) extracts domain-invariant features for pose estimation, while INDG-Fi (Yang et al., 2024b) and freeGait (Yan et al., 2024c) utilize domain randomization and adversarial techniques to ensure robust alignment. CATFSID (Wei et al., 2024), ADA (Zinys et al., 2021), and DAFI (Li et al., 2021c) combine domain classifiers with techniques like triplet losses and gesture discriminators to align features and improve cross-domain generalization. Furthermore, Shi et al. (Shi et al., 2020a) and DATTA (Strohmayer et al., 2024) focus on dynamic alignment using domain classifiers for test-time adaptation. Other works such as WiSR (Liu et al., 2023a), WiCAR (Wang et al., 2019d), and Berlo et al. (Berlo et al., 2023) explore multi-domain classifiers for cross-device and cross-environment sensing, while WiCross (Qin et al., 2023) uses classifiers for multi-environment adaptation.
Domain Alignment with Similarity Matrix employs similarity metrics such as Maximum Mean Discrepancy (MMD) or cosine similarity for feature alignment. Li et al. (Li et al., 2021e), Hassan et al. (Hassan et al., 2024) and WiSDA (Jiao et al., 2024) leverage cosine similarity to align source and target features, while LAGER (Chen et al., 2024a) and WiCAU (Cui et al., 2024) apply MMD and Wasserstein distance for distribution alignment. CrossFi (Zhao et al., 2024a) uses Siamese networks and MK-MMD to align distributions for few-shot and zero-shot adaptation. P3ID (He et al., 2024) combines DTW-based soft labeling with MMD, and AdapLoc (Zhou et al., 2020) aligns probability distributions using Euclidean distances. Other methods, such as WiAi-ID (Liang et al., 2023), Chen et al. (Mehryar, 2023) , and Mini-Batch Alignment (van Berlo et al., 2023, 2022), use advanced similarity metrics for fine-grained alignment. AirFi (Wang et al., 2022d) aligns multiple source domains using MMD and then applies it to the target domain by extracting common features from various training environments and minimizing the distribution differences between them. An arbitrary Gaussian noise is added to each CSI data sequence, generating new CSI datasets that help approximate the CSI data distribution in different environments, which benefits system training.
Domain Alignment with GAN focuses on generating synthetic samples for better alignment. MetaGanFi (Zhang et al., 2022b) integrates CycleGAN for domain alignment while leveraging meta-learning to enhance generalization.CsiGAN (Xiao et al., 2019) uses cyclegan (Zhu et al., 2017a), SIDA (Zhang et al., 2023b) uses VAE and GANto generate domain-specific data and align minority domains to the majority. Zhang et al. (Zhang et al., 2022a) uses CycleGAN to convert source features into target domain representations for improved cross-domain performance. WiLISensing (Ding et al., 2020b) first uses MMD to align multiple source domains and then leverages transfer learning to transfer features between different locations, freezing layers before the fully connected layers and training with a small number of samples to enhance model adaptability. PIAS (Xiao et al., 2024b) designs a cluster to generate pseudo-labels for target domain features and filters out unreliable training samples, using Contrastive Domain Discrepancy (CDD) loss to minimize intra-class feature discrepancies and maximize inter-class feature discrepancies. SCTRFL (Zheng et al., 2020) utilizes Restricted Boltzmann Machines (RBM) for amplitude and phase reconstruction in new environments. Distorted CSI amplitude and processed phase are input into trained autoencoders, which rebuild calibrated CSI from real-time measurements, restoring them to their reference state. Component separation with a discriminator is used in works like Elujide et al. (Elujide et al., 2022), which applies unsupervised adversarial invariance, and Person-in-WiFi (Wang et al., 2019f). UH-Sense (Chen et al., 2024b) uses GAN to remove noise from features, while Wi-Dist (Zhang et al., 2024d) uses CycleGAN to generate more data, training the classifier and feature generator adversarially to mitigate the negative impact of environmental changes. Wi-Piga (Hao et al., 2021) addresses personnel independence in action recognition using GANs, while Wang et al. (Wang et al., 2024d) decouples features related to identity and activity with separate task classifiers. In the domain-specific memory approach, CrossID (Wu et al., 2021a) learns affine parameters and identity attributes modeled as the interrelationship of affine parameters, which can be conditioned through instance normalization.
2.3.3. Meta-learning
Meta-learning techniques are designed to enable Wi-Fi sensing models to quickly adapt to new domains or tasks with minimal labeled data. These approaches are especially valuable when dealing with scenarios where training data is scarce or environments are highly dynamic. In the case of MetaFormer (Sheng et al., 2024), meta-learning is applied through the Prototypical Network (Snell et al., 2017), which learns class prototypes for few-shot classification tasks. MetaGanFi (Zhang et al., 2022b) uses CycleGAN to align multiple domains into a unified domain and integrates MAML (Finn et al., 2017) for efficient meta-learning, allowing the model to quickly adapt to new domains. WiLiMetaSensing (Ding et al., 2021a) adopts a matching network to learn domain-invariant representations by matching similar data points from different domains. Wi-AM (Zhang et al., 2023j) uses meta-learning techniques to optimize activity recognition, although the specific meta-learning method is not detailed. AFSL-HAR (Wang et al., 2021b) employs MAML (Finn et al., 2017) to adapt human activity recognition models across domains with minimal labeled data. RoMF (Zhang et al., 2024b) combines meta-learning with Graph Neural Networks (GNNs) for cross-domain activity recognition. The specific meta-learning algorithm is not identified, but the combination of graph-based learning with meta-learning enhances model performance across environments. Huang et al. (Huang et al., 2022) uses MAML (Finn et al., 2017) for efficient adaptation in localization tasks, enabling fast adaptation to new environments. Metaloc (Gao et al., 2023) uses MAML (Finn et al., 2017) to adapt localization models in new environments with minimal data, improving generalization to unseen domains. Zhang et al. (Zhang et al., 2020b) propose a meta-learning approach that enables models to be fine-tuned with very little training effort in new environments or for recognizing new types of activities, based on the Learning to Learn principle (Andrychowicz et al., 2016). In TOSS (Zhou et al., 2022b), a dynamic pseudo-labeling strategy is combined with MAML for few-shot learning. The model uses both labeled and unlabeled target samples, leveraging MAML to adapt quickly to new domains. ML-WiGR (Gao et al., 2022b) uses Reptile (Nichol, 2018), another meta-learning algorithm, to improve the generalization of Wi-Fi sensing systems across different domains. Similarly, Wi-learner (Feng et al., 2022) employs MAML and Doppler Frequency Shift (DFS) for rapid domain adaptation in human activity recognition tasks. Finally, RF-Net (Ding et al., 2020a) uses metric-based meta-learning for one-shot human activity recognition, enabling the model to recognize new activities with only a single labeled example from each class.
2.3.4. Embedding Learning
Embedding learning focus on constructing compact, robust, and domain-invariant feature representations to enhance the adaptability of Wi-Fi sensing systems in diverse domains. These methods are especially useful for scenarios like few-shot learning, zero-shot learning, and cross-domain recognition. Some approaches leverage few-shot and prototypical learning frameworks, such as FewCS (Yang et al., 2023b) and ReWiS (Bahadori et al., 2022), which utilize prototypical networks to learn class prototypes for robust cross-domain recognition. Similarly, WiGesID  (Zhang et al., 2022c) employs relation networks to model the relationships between input samples and improve the generalization of domain-invariant embeddings, while MatNet-eCSI (Shi et al., 2020b) integrates matching networks for effective few-shot learning in varying domains. LT-WiOB (Zhou et al., 2022c) and TransferSense  (Bu et al., 2022) use triplet losses to generate embeddings that preserve inter-class separation and intra-class compactness, enabling better performance in unseen environments. For re-identification and activity recognition, 3D-ID  (Ren and Yang, 2022) uses 2D AoA estimations and Siamese networks to construct embeddings for 3D body representations, enabling robust person re-identification across domains. Similarly, WiDeo (Islam and Nirjon, 2021) combines Wi-Fi and visual data to generate joint embeddings that enhance activity recognition across domains, while CrossFi  (Zhao et al., 2024a) integrates Siamese networks with MK-MMD for zero-shot and few-shot domain adaptation. Other embedding learning approaches rely on graph-based learning and pretext tasks. SFTSeg  (Xiao et al., 2022) applies time-series-specific pretext tasks and triplet loss to train models with limited labeled samples, while CSI-GDAM (Zhang et al., 2021a) leverages Graph Neural Networks (GNNs) and relation networks to create domain-invariant activity representations.
Some methods focus on integrating multi-view or triplet embeddings. For example, UniFi (Liu et al., 2024b) employs spatio-temporal multi-view networks with attention mechanisms and triplet loss to align multi-device embeddings, while MaP-SGAN (Liang et al., 2024) uses GANs to augment data and train Siamese networks for robust cross-domain embeddings. Similarly, Wi-Cro (Mao et al., 2024) combines triplet loss with synthetic data generated by CycleGAN for cross-domain adaptation. Loss function-based embedding approaches, such as CATFSID (Wei et al., 2024), WiOpen (Zhang et al., 2024c), and WiGR (Hu et al., 2021a), also utilize triplet losses to align source and target domain embeddings, ensuring robustness against domain-specific noise. In the context of matching networks and meta-learning, DAFI (Li et al., 2021c) uses matching networks with triplet losses to enhance cross-domain feature alignment. Other methods, like Chen et al. (Chen and Chang, 2022), Ding et al. (Ding et al., 2021b), and WiONE (Gu et al., 2021), apply matching and prototypical networks for cross-domain authentication and activity recognition, while Shi et al. (Shi et al., 2022), Caution (Wang et al., 2022e), and LT-WiOB (Zhou et al., 2022d) improve few-shot generalization through similar frameworks. GAN-based methods like Wang et al. (Wang et al., 2020) and Zhou et al. (Zhao et al., 2024c) integrate GAN-generated samples into prototypical embeddings to improve cross-domain recognition performance. Furthermore, methods such as CSI-GLSTN (Zhang et al., 2024e) use GNNs to learn hidden activity graph representations, enabling robust cross-domain activity detection, while HAR-MN-EFF (Shi et al., 2020c) and DA-HAR (Sheng et al., 2023) enhance embeddings for human activity recognition through matching networks and triplet loss.
[…]

### v2

@@ II-C Feature Learning Stage
The Feature Learning Stage is fundamental to improving the adaptability and robustness of Wi-Fi sensing systems across diverse domains. This stage includes several key techniques aimed at overcoming the challenges of domain variability, limited labeled data, and environmental changes. Methods such as domain alignment, metric learning, data augmentation, and pseudo-labeling work together to enhance model performance, ensuring that features learned from one domain can be successfully applied to others. These approaches through techniques like adversarial training, similarity metrics, data synthesis, and embedding refinement, enable Wi-Fi sensing models to generalize effectively in dynamic and heterogeneous real-world environments.
II-C1 Domain Alignment
Adversarial Domain Adaptation [241, 248] aims to align the source and target domains, enabling Wi-Fi sensing to generalize across new individuals, environments, and other novel domains.
Domain Alignment with Domain Discriminator
Figure 7: Domain alignment with domain discriminator or domain classifier.
Fig. 7 illustrates a representative domain-adversarial training framework for Wi-Fi sensing generalization. The model takes both source-domain and target-domain Wi-Fi data as input and employs a deep neural network to extract shared features. These features are then fed into two branches: an action classifier for activity recognition, and a domain discriminator that attempts to determine whether the input originates from the source or target domain. Through adversarial training, the feature extractor learns to produce domain-invariant representations that are disentangled along the domain dimension, thereby enabling more robust cross-domain activity recognition. This fundamental framework was employed in WiHARAN [117], DA-HAR [118], and Zhang et al. [119] to improve cross-environment activity recognition.
Several variants have been proposed to enhance this domain-adversarial learning framework further. For example, i-Sample [120] adopts a two-stage training strategy. In the first stage, the model focuses on learning a domain-invariant feature extractor by jointly feeding source and target domain data. A domain discriminator is used to classify the domain origin of each sample, while a parallel branch performs activity classification. In the second stage, the feature extractor is frozen, and the classifier is fine-tuned using labeled source data. This staged approach effectively enhances cross-environment generalization performance. In another line of work, JADA [121] employs two separate encoders: one for the source domain and another for the target domain. Through adversarial training with a domain discriminator, the features extracted from both encoders are encouraged to become domain-invariant. After training, the source-domain encoder is frozen, and a classification head is trained using labeled source data. The final model—comprising the target-domain encoder and the trained classification head—is used for inference on target-domain samples, achieving robust cross-domain activity recognition. Other examples are CrossGR [122] and Wi-Adaptor [123], which replace the traditional domain discriminator with an auxiliary action classifier. Unlike standard classifiers, this branch is trained to assign uniform probabilities across all action classes, effectively making actions indistinguishable in the feature space. This strategy forces the feature extractor to remove task-relevant cues from the domain-variant features, thereby encouraging it to learn more domain-invariant and task-discriminative representations. As a result, the main action classifier benefits from cleaner, more generalizable features during training.
Domain Alignment with Domain Classifier
Another line of work extends the domain-adversarial framework by replacing the domain discriminator with a domain classifier. While the standard domain discriminator performs binary classification to distinguish between source and target domains, the domain classifier generalizes this concept to multi-class classification, where each domain (e.g., different environments or deployment scenarios) is assigned a distinct label, also shown in Fig. 7. For example, EI [124] introduces a domain classifier to explicitly recognize the environment in which the activities are recorded. Through adversarial training, the domain classifier is optimized to correctly classify each environment, while the feature extractor is trained to fool the domain classifier, thereby learning environment-invariant representations. Similarly, Khattak and Khan [125] and WiCAR [126] adopt this strategy to achieve location-invariant activity recognition and in-car environment-invariant activity recognition, respectively.
Beyond environments and locations, user identity can also serve as a domain label. Systems such as GESFI [127], ADA [128], and DATTA [129] use the user ID as the domain label to enable user-independent gesture recognition. This idea can be extended further: domain labels include user identity, body orientation, physical position, and environment. Building on this fine-grained decomposition, models like WiSR [130], Yin et al. [93], INDG-Fi [131], WiCross [64], and Berlo et al. [132] demonstrate the feasibility of learning representations that generalize across multiple domain factors simultaneously.
It is worth noting that this domain-classifier-based adversarial framework is not limited to activity recognition. By replacing the action classifier with other task-specific heads, this architecture has been successfully applied to a variety of tasks, including: pose estimation across subjects (e.g., DINN [133]), gait recognition across actions (e.g., freeGait [134]), indoor localization across environments (e.g., DAFI [135]), and user identification across locations (e.g., Shi et al. [106]). These extensions further underscore the versatility and effectiveness of domain-classifier-based adversarial training in improving generalization across a wide range of Wi-Fi sensing applications. In addition, Fidora [136] replaces the domain classification branch with a data reconstruction branch. In this framework, both source and target domain samples are first encoded by a shared feature extractor, and then passed through a decoder to reconstruct the original input data. The reconstruction loss serves as an indirect objective for domain alignment, encouraging the feature extractor to retain domain-invariant information that is useful for reconstruction.
Domain Alignment with Similarity Computing
Figure 8: Maximum Mean Discrepancy [137] quantifies the similarity between two probability distributions by projecting data into a Reproducing Kernel Hilbert Space (RKHS) and computing the squared distance between their mean embeddings, encouraging the model to learn domain-invariant representations.
Another widely adopted strategy for domain alignment is to explicitly minimize the distribution discrepancy between source and target domains. One representative approach is Maximum Mean Discrepancy (MMD) [137], which quantifies the difference between two probability distributions by projecting data into a Reproducing Kernel Hilbert Space (RKHS) and computing the squared distance between their mean embeddings, encouraging the model to learn domain-invariant representations, as shown in Fig. 8. In the context of Wi-Fi sensing, MMD has been successfully integrated into model objectives to mitigate domain shifts and enhance cross-domain generalization. For example, AdaPose [138] and BSWCLoc [139] utilize MMD-based alignment for cross-environment human pose estimation and indoor localization, respectively.
Beyond standalone MMD alignment, some works combine MMD with domain classifiers to further enhance generalization. For instance, Kang et al. [140] introduce a dual-branch architecture, where MMD is applied between the source and target domains at the output of the activity classifier, while a parallel domain classifier is trained to align features across domains. This design enables robust activity recognition across variations in user identity, environment, location, and body orientation. Similarly, WiAi-ID [141] adopts MMD to align source and target features, while employing a domain classifier to discriminate appearance characteristics, achieving appearance-independent passive person identification.
In addition to MMD, several alternative metrics have been adopted to quantify distributional similarity between source and target domains. For example, Wi-CHARS [142] applies Earth Mover’s Distance (EMD) to align features for cross-environment activity recognition. Mehryar [143] utilizes the Kullback–Leibler (KL) divergence to reduce domain shift, while AdapLoc [144] employs Euclidean distance for indoor localization across domains. PIAS [145] further explores Contrastive Domain Discrepancy (CDD) [146] to align representations in a contrastive manner. WiCAU [147] uses Wasserstein distance to measure to quantify distributional similarity between source and target domains.
In some cases, in addition to knowing the domain labels (e.g., environment), the class labels (e.g., action types) of both source and target domain samples are also available during training. This enables a finer-grained alignment approach, where domain-level similarity measures can be replaced by class-level distribution alignment, allowing for more precise cross-domain representation learning. A number of studies have explored this strategy under different names and formulations. For example, local MMD [249, 212], mini-batch alignment [98, 99], Multiple Kernel MMD [179], and sub-domain alignment [250] all share the common idea of aligning corresponding action classes across domains, instead of aligning the global distributions. The similarity metrics used for sub-domain alignment vary across works. For instance, P3ID utilizes Euclidean distance to measure the closeness between class features from different environments. WiSDA [148], FDAS [149], and CDFi [150] adopt cosine similarity to measure the closeness between class-conditional features from different domains. Notably, CDFi further incorporates Jensen–Shannon (JS) divergence [251] to perform an additional global domain alignment, complementing the sub-domain matching process.
In cases where only domain-level labels (e.g., environment, location) are available, but class labels of the target domain samples are unknown, some approaches generate pseudo-labels for target domain data during training to enable the above fine-grained domain alignment. For example, LAGER [96] and Zhan & Wu [151] adopt this strategy by first predicting class labels for the unlabeled target domain samples. These pseudo-labeled data are then used to perform sub-domain alignment. Both methods use Euclidean distance to measure the similarity between class-conditional feature distributions. LAGER achieves domain generalization in cross-user, cross-person, cross-location, and cross-orientation activity recognition, while Zhan & Wu’s approach targets cross-scene indoor localization.
When only multiple source domains are available during training and no target domain data is accessible, models can still improve generalization by aligning the source domains themselves. For example, AirFi [152] performs MMD-based alignment across multiple source domains, learning domain-invariant features that generalize to unseen target environments for activity recognition. Similarly, WiLISensing [153] adopts an MMD alignment strategy across source domains to train the feature extractor. During deployment, the feature extractor is frozen, and only a small number of labeled samples from the new environment are used to train the final fully-connected layers. This approach effectively enables cross-environment activity recognition with minimal adaptation effort.
Domain Alignment with Generative Adversarial Networks
Figure 9: Generative Adversarial Networks
Another line of research explores Generative Adversarial Networks (GANs) [252] to perform domain alignment by transforming data distributions or learning shared latent representations. GAN-based alignment builds on the adversarial learning mechanism of GANs, shown in Fig. 9, which involve a generator and a discriminator engaged in a minimax game. The generator aims to produce data indistinguishable from real samples, while the discriminator attempts to differentiate between real and generated data. Through iterative optimization, the generator learns to synthesize data that closely follows the expected distribution. For example, MetaGanFi [154] proposes a three-stage pipeline: (1) In the first stage, a CycleGAN-based framework [253] is used to convert Wi-Fi data from multiple source domains into a unified Uni-domain representation. This is achieved through adversarial training, where the generator transforms source-domain data to resemble a common distribution, and a discriminator is trained to identify the originating domain. The generator is optimized to fool the discriminator, thereby removing domain-specific cues. (2) In the second stage, the generated Uni-domain data are used to train an activity classifier, ensuring that the classifier learns from domain-invariant representations. (3) Finally, during inference, target-domain data are passed through the trained generator and classifier to perform activity recognition for unseen users. Similarly, Zhang et al. [155] and Wi-Cro [156] and WiTeacher [157] employ CycleGAN and StyleGAN [254], respectively, but instead of mapping source domains to a unified representation, they transform source domain data to match the target domain distribution, thereby enabling cross-domain sign language recognition and activity recognition.
Domain Alignment with Multi-task Learning
Multi-task learning improves generalization by jointly training a model on multiple related tasks or domains, either through training multiple expert networks or by sharing parameters across tasks. This approach encourages the learning of representations that are both domain-invariant and task-relevant.
For example, CrossSense [158] trains 10 expert models. The experts are trained offline, and at runtime, the appropriate expert for a given input is automatically chosen. This expert selection mechanism allows the system to adaptively handle different input conditions, improving robustness across domains. Similarly, Sugimoto et al. [159] propose a system that employs a multi-task learning approach based on an encoder-decoder network architecture, where each decoder corresponds to a specific target environment. This structure ensures that the encoder learns a more general representation that can accommodate the diversity between different environments. CSI-MTGN [160] also adopts multi-task learning to learn shared parameters for the target task. Instead of using separate decoders or experts, CSI-MTGN builds a unified model where parameters are shared across multiple tasks. This shared representation facilitates knowledge transfer between tasks and domains, enhancing the model’s ability to generalize from limited data. By integrating shared learning modules and task-specific heads, CSI-MTGN effectively balances generalization and specialization, making it well-suited for location-aware sensing tasks with high domain variability.
Domain Alignment with Cross-modal Embedding
Cross-modal alignment leverages the rich semantic information inherent in modalities such as text, images, and radar signals, which are often better understood by large-scale pre-trained models (e.g., language models or vision-language models) than raw Wi-Fi CSI. By mapping Wi-Fi data into these semantically meaningful spaces, models can benefit from external knowledge and enhanced generalization. Consequently, several works explore cross-modal embedding techniques to bridge CSI with modalities such as text, vision, and mmWave radar, enabling more robust domain adaptation.
Wi-Fringe [161] converts action class labels into BERT embeddings [255] and uses them as targets to guide the learning of Wi-Fi representations. This enables the system to bridge the semantic gap between textual activity descriptions and Wi-Fi signal patterns. Such alignment with language models facilitates better generalization to unseen activities, particularly in open-set scenarios where the label space at test time differs from training. XRF55 [70] also employs BERT embeddings as additional supervision to enhance the semantic alignment of Wi-Fi feature learning. Wi-Chat [162] goes one step further by using natural language to describe Wi-Fi signal patterns corresponding to behaviors such as walking, falling, no-event, and breathing. These descriptions are provided to a large language model (GPT-4o), enabling it to understand and interpret Wi-Fi sensing data. This approach enables zero-shot action recognition and respiratory monitoring, demonstrating a novel integration of wireless sensing with LLMs.
In addition to text-based alignment, radar signals also offer fine-grained motion information that complements Wi-Fi data. WiFi2Radar [163] proposes a cross-modal translation framework that converts CSI spectrograms into mmWave radar Doppler spectrograms using a U-Net [256] trained with synchronized radar data. This allows the model to inherit radar’s high-resolution motion sensitivity while retaining the deployment convenience of Wi-Fi infrastructure.
II-C2 Component Disentangle
Wi-Fi signals inherently contain entangled information about the environment, human presence, and motion. Several works aim to disentangle these components to enable cross-domain generalization.
For example, Elujide et al. [164] leverage unsupervised adversarial invariance [257] to disentangle gesture-related features from location-related features in the latent space, thereby enabling gesture recognition in unseen locations. UH-Sense [165] employs a GAN-based framework to denoise the data affected by UAV jitter. By filtering out motion-induced distortions, the model achieves robust human localization based on Wi-Fi signals collected from UAV platforms. Person-in-WiFi [45] assumes that Wi-Fi signals are composed of both environment-related and motion-related components. It uses a GAN to transform the signal representation into an environment-invariant latent space—one in which a domain discriminator cannot predict the environment index—thereby improving generalization for cross-environment human pose estimation. Wi-Piga [166] assumes that Wi-Fi signals contain identity-related and action-related components. It utilizes a GAN to generate features in which the identity information is no longer distinguishable. These identity-invariant features are then used for yoga pose recognition, enabling cross-user generalization. CrossID [167] leverages conditional instance normalization [258] to learn affine parameters that capture identity attributes. During inference, these parameters can be adapted to new users, enabling gesture recognition for previously unseen individuals.
II-C3 Metric Learning for Zero/Few-shot Learning
Figure 10: Triplet loss aims to structure the feature space such that samples from the same class are mapped close together ($x_{1}^{+},x_{2}^{+}$), while samples from different classes are pushed farther apart ($x^{+},x^{-}$).
Metric learning is another important strategy for Wi-Fi sensing generalization, which aims to structure the feature space such that samples from the same class are mapped close together ($x_{1}^{+},x_{2}^{+}$), while samples from different classes are pushed farther apart ($x^{+},x^{-}$). This is typically achieved during training by optimizing pairwise or triplet losses over sample embeddings, as shown in Fig. 10. After training, the resulting feature space forms well-separated clusters corresponding to different semantic classes.
At inference time, unlike traditional classification approaches that directly predict class labels, metric learning relies on a support set, which contains representative examples (templates) for each class. For example, in cross-environment activity recognition, the support set may consist of one labeled sample per activity from the target environment. To classify a query sample, its feature is extracted via the trained feature encoder and compared against the features of all support set samples using a similarity metric such as cosine or Euclidean distance. The class label of the closest support sample is then assigned to the query. This enables few-shot inference in the target domain with only a handful of labeled examples. Furthermore, if the support set is constructed using only training data without any target-domain samples, the model can perform zero-shot inference, which makes it particularly suitable for domain generalization scenarios.
Metric Learning with Triplet Loss
Several works have leveraged triplet loss, as illustrated in Fig. 10, for cross-domain few-shot Wi-Fi sensing. For example, LT-WiOB [168] and CATFSID [169] adopt the classic triplet loss strategy to train feature extractors that pull embeddings of intra-class samples closer while pushing those of inter-class samples further apart. During inference, a support set is constructed using a few labeled samples per class from the target domain. The test sample is then classified by computing distances to the support set embeddings and assigning the label of the nearest neighbor, thereby enabling cross-environment activity recognition and user identification. Similarly, TransferSense [170] employs triplet loss for cross-domain gait recognition, where the support set consists of a few samples from previously unseen users in a new environment, enabling few-shot identification based on walking patterns.
While most approaches use the standard Euclidean distance to compare embeddings, recent methods explore alternative similarity measures to improve robustness. UniFi [171] introduces a mutual information-based loss [259] to enhance discrimination between positive and negative pairs, facilitating cross-scene, cross-location, and cross-orientation activity recognition. WiOpen [172] adopts a metric learning strategy based on neighbor component analysis (NCA) [173], which minimizes intra-class distances while maximizing inter-class separability, thereby improving few-shot classification on unseen gesture classes.
Beyond handling domain and user shifts, triplet-loss-based models have also been applied to cross-modality recognition. For instance, SFTSeg [174] trains a model on the IMU-based HandGesture dataset [260] and transfers it to the Wi-Fi-based WiFiAction dataset [261], achieving few-shot recognition across sensing modalities.
Learning with Contrastive Learning
Contrastive learning also relies on constructing positive and negative sample pairs, and often incorporates data augmentation to expand the training set. Unlike triplet loss, where each training iteration involves a single positive and negative pair, contrastive learning considers all pairwise relationships within a mini-batch. Specifically, for a mini-batch of $n$ samples, an $n\times n$ label matrix $X={x_{i,j}}$ is constructed, where each element $x_{i,j}$ indicates whether sample $i$ and sample $j$ belong to the same class (1) or not (0). The diagonal elements are always 1, as each sample matches itself. The loss is typically computed as a row-wise or column-wise cross-entropy over this similarity matrix, enabling the model to learn class-discriminative embeddings across all pairs in the batch. During inference, a query sample is compared with support samples in the feature space, and the label of the most similar support sample is assigned as the prediction result.
Several recent works have demonstrated the effectiveness of contrastive learning for few-shot and cross-domain Wi-Fi sensing tasks. DualConFi[175] applies a contrastive learning framework to enable robust activity recognition across different environments. CFLH [176] enhances the diversity of positive sample pairs by applying random scaling augmentations to Wi-Fi signals, and then trains the model using contrastive loss to enable few-shot recognition of unseen activities in new environments. Xiao et al. [177] further integrate diffusion-based generative modeling (e.g., Denoising Diffusion Probabilistic Models [262]) to synthesize additional positive samples. These synthetic samples are incorporated into a contrastive training pipeline, which improves the model’s ability to recognize novel activities in unseen environments using limited labeled data.
Metric Learning with Siamese Neural Networks
Figure 11: Siamese Neural Networks
Siamese neural networks [263] offer a classical yet effective architecture for few-shot learning and domain generalization. In this framework, as shown in Fig. 11, a pair of samples is fed into two identical networks that share weights. Each network extracts feature embeddings from its input, and the resulting embeddings are concatenated and passed through a binary classifier. The classifier is trained to predict whether the two inputs belong to the same class: if they do (e.g., two samples of the same activity), the target label is 1; otherwise, it is 0. During inference, a support set is constructed from a few labeled samples per class in the target domain. To classify a query sample, it is paired with each support sample and processed by the Siamese network. The network outputs a confidence score indicating the likelihood of the two samples belonging to the same class. The class label of the support sample with the highest confidence score is then assigned to the query sample.
This pipeline has been widely adopted in Wi-Fi sensing tasks for few-shot cross-domain generalization. For instance, 3D-ID [116] utilizes this pipeline for few-shot user identification in unseen environments. MaP-SGAN [178] applies it for gait recognition by enhancing the diversity of training samples via generative models. CrossFi [179] extends the Siamese framework to handle both cross-domain and new-class scenarios, enabling gesture recognition and user identification in previously unseen domain.
Metric Learning with Relation Network
Figure 12: Relation Network
Relation networks [264] provide a powerful framework for few-shot learning by explicitly modeling the similarity between samples through deep relational reasoning. As illustrated in Fig. 12, training involves two components: a support set and a query set. The support set contains one labeled sample per class, while the query set contains a sample whose class is to be predicted. Both support and query samples are passed through a shared feature extractor. The resulting feature vectors are then concatenated and passed to a relation module (typically a neural network), which outputs a similarity score indicating how likely the two samples belong to the same class. The network is optimized using a multi-class cross-entropy loss, where the ground-truth label is encoded in a one-hot format: for each query sample, only the score corresponding to the matching support class is set to 1, while all others are 0. During inference, a test support set and a query sample are fed into the model. The network outputs relation scores between the query and each support sample. The class corresponding to the support sample with the highest relation score is then selected as the prediction result.
Wang et al. [180] follow the above pipeline to enable few-shot activity recognition in unseen environments. WiGR [181] adopts relation networks and trains the model on 100 action classes. To evaluate generalization to novel classes, the model is tested on previously unseen activities, using only three support samples per new class. LESS [182] also uses relation networks ito achieve few-shot cross-environment fingerprint-based localization. WiGesID [183] further evaluates this framework to simultaneously support cross-environment action recognition and user identification. Besides, WiGesID also supports few-shot classification of both new actions and unseen individuals, showcasing the flexibility of relation networks in multi-task and cross-domain scenarios.
Metric Learning with Matching Network
Figure 13: Matching Network
Matching networks [265] provide an effective framework for few-shot learning by leveraging attention mechanisms and episodic training to mimic test-time inference. As illustrated in Fig. 13, during training, each episode includes a support set $S=\{(x_{i},y_{i})|i=1,2,...,n\}$, where $x_{i}$ is a sample, $y_{i}$ is its corresponding label, and $q$ is a query sample. The support set consists of a few labeled samples from each class, while the query sample is to be classified.
The predicted label of the query sample is computed as a weighted sum of the support set labels, where the weights are derived from the similarity between the query and each support sample:
$a(q,x_{i})$ $={e^{c(f(q),g(x_{i}))}}/{\sum_{j=1}^{n}e^{c(f(q),g(x_{j}))}}$ (1)
[…]